# Rethinking Attention: Polynomial Alternatives to Softmax in Transformers

### v2

@@ A.1 Theoretical analysis
$||\mathbf{softmax}(A)||_{F}^{2}$ $=\bigg{(}\frac{1}{\sum_{j=1}^{n}e^{a_{1j}}}\bigg{)}^{2}(e^{2a_{11}}+\cdots e^{2a_{11}})+\cdots$ (13)
@@
It then immediately follows that $||\mathbf{softmax}(A)||_{F}\leq\sqrt{N}$ and this proves the first inequality in the statement of theorem 4.1.
@@
$||\nabla\mathbf{softmax}(A)||_{F}^{2}=||\nabla F_{11}||_{F}^{2}+\cdots+||\nabla F_{NN}||_{F}^{2}.$ (25)
We will proceed by bounding each collection $||\nabla F_{i1}||_{F}^{2}+\cdots+||\nabla F_{1N}||_{F}^{2}$ separately then add up all the bounds. We have
$||\nabla F_{i1}||_{F}^{2}+\cdots+||\nabla F_{1N}||_{F}^{2}$ $=|F_{i1}-F_{i1}^{2}|^{2}+|F_{i1}F_{i2}|^{2}+\cdots+|F_{i1}F_{iN}|^{2}$ (26)
@@
$\hskip 14.22636pt+4F_{iN}^{2}(F_{i1}+F_{i2}+\cdots+\widehat{F_{iN}})$ (36)
@@
Putting all the bounds together for each of the terms $N$ terms $||\nabla F_{i1}||_{F}^{2}+\cdots+||\nabla F_{1N}||_{F}^{2}$ we get
$||\nabla\mathbf{softmax}(A)||_{F}^{2}\leq 4N$ (39)
@@
$||\nabla\mathbf{softmax}(A)||_{F}\leq 2\sqrt{N}.$ (40)
@@
In this section we will give the proof of theorems 4.2 and 4.4.
@@
For the case that $p>1$ we proceed in a similar way. The key observation is that odd powers, in the matrix expansion, will have expectaion 00, so we need only consider the even powers. Therefore, suppose $\mathbf{C}=(\mathbf{A}\mathbf{B})^{p}$. We will compute the expectation of the first entry $c_{1}1\in\mathbf{C}$:
@@
Note that the first term only has a count of $D$ and the second term has a count of $D(D-1)$. Thus, we only need to consider the $\mathcal{O}(D^{p})$ term where all the components have a power of 2. The count is similar to choosing $p$ items from $D$,
@@
$=\frac{D!}{(D-p)!}(2p-1)!!\sigma_{1}^{2p}\sigma_{2}^{2p}.$
@@ A.2.1 Hardware
The vision transformer experiments in Section 6.1, the object detection and instance segmentation experiments in Section 6.2 and the Nyströmformer experiments from Section 6.3 were all carried out on Nvidia A100 GPUs. The physics informed transformer experiments in Section A.2.5 were carried out on a Nvidia A6000 GPU.
@@ A.2.2 Experimental hyperparameters
Vision transformers in Section 6.1.
In Section 6.1 we tested four different vision transformers, ViT-B [10], DeiT-B [25], Swin-B [17] and XCiT-M [3], with the activations softmax, $x^{3}$ + fixed scaling, $x^{3}$ + dynamic scaling and $x^{3}$. The training strategy follow the exact strategy used in each of the original papers, we used the Timm [28] libraries to train our models.
Object detection and instance segmentation in Section 6.2.
@@
Nyströmformer in Section 6.3.
@@ A.2.3 Frobenius norm computations
In Section 5 we showed plots of the Frobenius norm of the self-attention matrix and for the Jacobian of the self-attention matrix for softmax, $\frac{1}{14}x^{3}$, and $x^{3}$. This was done for a ViT-Tiny architecture on the Tiny-ImageNet dataset. Figure 6 shows the plots of the Frobenius norm of the self-attention matrix for the ViT-Tiny architecture, during training, for all layers averaged over the heads within each layer. Figure 7 shows the Frobenius norm of the Jacobian of the self-attention matrix during training for each layer, averaged over the total number of heads within each layer.
@@ A.2.4 Discussion on Linear attention and other polynomials on ViTs
In the experiments Section 6 we showed our insights using the polynomial $x^{3}$ as this was a non-linear polynomial that did not satisfy the key properties that softmax did. In this section we will compare softmax with other polynomials that we also found to perform well.
The results for ViTs is shown in Table 5. As we can see from that table each polynomial does far better when scaled using the theory from Section 4. We note that when we tried to train polynomial higher than order $6$ they did not train well. On inspecting the weights we found that several were very close to zero leading to a gradient vanishing problem. We hypothesize that this is because the transformers were randomly initialized with weights about zero. Thus taking such a weight and applying a polynomial of the form $x^{k}$ with $k\geq 6$ would make those weights orders of magnitude smaller, making it difficult to train after some point.
@@ A.2.5 Physics informed transformers
As in [31] we considered a physics informed transformer (PINNsformer) on the above two PDEs. We compared softmax with $x^{3}$ + fixed scaling, $x^{3}$ + dynamic scaling, where the scale was initialized to $\frac{1}{\sqrt{51}}$ as the input sequence length is $51$, and $x^{3}$. As in [31] we computed the training loss, and for testing the relative mean absolute error (rMAE), and the relative mean squared error (rMSE). Table 6 shows the results from which it can be seen that both $x^{3}$ + fixed scaling and $x^{3}$ + dynamic scaling are competitive with the original softmax.

### v3

@@ 9 Conclusion
10 Acknowledgements
Simon Lucey acknowledges support from the Australian Research Council (ARC) through the Discovery Project DP220103803.
@@ A.1 Theoretical analysis
[…]
@@
It then immediately follows that $\mathbf{softmax}(A)||_{F}\leq\sqrt{N}$ and this proves the first inequality in the statement of theorem 4.1.
@@
[…]
@@
$\hskip 14.22636pt+\cdots\cdots\cdots\cdots+$ (35)
[…]
@@
[…]
$\nabla\mathbf{softmax}(A)||_{F}^{2}\leq 4N$ (39)
@@
$\nabla\mathbf{softmax}(A)||_{F}\leq 2\sqrt{N}.$ (40)
@@
In this section we will give the proof of theorems 4.2 and $4.$
@@
For the case that $p>1$ we proceed in a similar way. The key observation is that odd powers, in the matrix expansion, will have expectaion 0, so we need only consider the even powers. Therefore, suppose $\mathbf{C}=(\mathbf{A}\mathbf{B})^{p}$. We will compute the expectation of the first entry $c_{1}1\in\mathbf{C}$:
@@
Note that the first term only has a count of $D$ and the second term has a count of D(D1)D(D-1). Thus, we only need to consider the $\mathcal{O}(D^{p})$ term where all the components have a power of 2. The count is similar to choosing $p$ items from $D$,
@@
$=\frac{D!}{(D-p)!}(2p-1)!\sigma_{1}^{2p}\sigma_{2}^{2p}.$
@@ A.2.1 Hardware
The vision transformer experiments in Section˜6.1, the object detection and instance segmentation experiments in Section˜6.2 and the Nyströmformer experiments from Section˜6.3 were all carried out on Nvidia A100 GPUs. The physics informed transformer experiments in Section˜A.2.5 were carried out on a Nvidia A6000 GPU.
@@ A.2.2 Experimental hyperparameters
Vision transformers in Section˜6.1.
In Section˜6.1 we tested four different vision transformers, ViT-B [10], DeiT-B [25], Swin-B [17] and XCiT-M [3], with the activations softmax, $x^{3}$ + fixed scaling, $x^{3}$ + dynamic scaling and $x^{3}$. The training strategy follow the exact strategy used in each of the original papers, we used the Timm [28] libraries to train our models.
Object detection and instance segmentation in Section˜6.2.
@@
Nyströmformer in Section˜6.3.
@@ A.2.3 Frobenius norm computations
In Section˜5 we showed plots of the Frobenius norm of the self-attention matrix and for the Jacobian of the self-attention matrix for softmax, $\frac{1}{14}x^{3}$, and $x^{3}$. This was done for a ViT-Tiny architecture on the Tiny-ImageNet dataset. Figure˜6 shows the plots of the Frobenius norm of the self-attention matrix for the ViT-Tiny architecture, during training, for all layers averaged over the heads within each layer. Figure˜7 shows the Frobenius norm of the Jacobian of the self-attention matrix during training for each layer, averaged over the total number of heads within each layer.
@@ A.2.4 Discussion on Linear attention and other polynomials on ViTs
In the experiments Section˜6 we showed our insights using the polynomial $x^{3}$ as this was a non-linear polynomial that did not satisfy the key properties that softmax did. In this section we will compare softmax with other polynomials that we also found to perform well.
The results for ViTs is shown in Table˜5. As we can see from that table each polynomial does far better when scaled using the theory from Section˜4. We note that when we tried to train polynomial higher than order $6$ they did not train well. On inspecting the weights we found that several were very close to zero leading to a gradient vanishing problem. We hypothesize that this is because the transformers were randomly initialized with weights about zero. Thus taking such a weight and applying a polynomial of the form $x^{k}$ with $k\geq 6$ would make those weights orders of magnitude smaller, making it difficult to train after some point.
@@ A.2.5 Physics informed transformers
As in [31] we considered a physics informed transformer (PINNsformer) on the above two PDEs. We compared softmax with $x^{3}$ + fixed scaling, $x^{3}$ + dynamic scaling, where the scale was initialized to $\frac{1}{\sqrt{51}}$ as the input sequence length is $51$, and $x^{3}$. As in [31] we computed the training loss, and for testing the relative mean absolute error (rMAE), and the relative mean squared error (rMSE). Table˜6 shows the results from which it can be seen that both $x^{3}$ + fixed scaling and $x^{3}$ + dynamic scaling are competitive with the original softmax.